# Fan distributions via Tverberg partitions and Gale duality

### v1

@@ Abstract.
Equipartition theory, beginning with the classical ham sandwich theorem, seeks the fair division of finite point sets in $\mathbb{R}^{d}$ by full-dimensional regions determined by a prescribed geometric dissection of $\mathbb{R}^{d}$. Here we examine equidistributions of finite point sets in $\mathbb{R}^{d}$ by prescribed low dimensional subsets. Our main result states that if $r\geq 3$ is a prime power, then for any $m$-coloring of a sufficiently small point set $X$ in $\mathbb{R}^{d}$ there exists a certain type of $r$-fan in $\mathbb{R}^{d}$ – namely, the union of $r$ “half-flats” of codimension $r-2$ centered about a common $(r-1)$-codimensional affine subspace – which captures all the points of $X$ in such a way that each half-flat contains at most an $r$-th of the points from each color class. The number of points in $\mathbb{R}^{d}$ we require for this is essentially tight when $m\geq 2$, and for typical point sets the interiors of the half-flats must share a robust fraction of the points of $X$. We also give analogous equidistribution results for small point sets in $\mathbb{C}^{d}$ by regular $r$-fans which are centered about a complex hyperplane.
In the same way that a theorem of Dolnikov on hyperplane piercings generalizes the ham sandwich theorem, our equidistribution results extend to “piercing distributions” for families of subsets of $X$ satisfying particular intersection criteria. Our proof method continues recent work of Frick et al., where the Gale transform was shown to link Dolnikov-type hyperplane transversals and Radon-type partitions. Similarly, our results arise as the real and complex Gale duals of linear versions of Tverberg-type theorems. Finally, we extend our distribution results to multiple $r$-fans after establishing a multiple intersection version of a topological Tverberg-type theorem due to Sarkaria.
@@ 1.1. From Equipartitions to Equidistributions of Colored Point Sets
Instead of equipartitions of finite point sets in $\mathbb{R}^{d}$ by full dimensional regions, here we seek their equidistribution by small-dimensional subsets of $\mathbb{R}^{d}$ arising from families of hyperplane with prescribed intersections. Specifically, suppose that $1\leq k\leq\min\{d-1,r-1\}$ and that $H_{1},\ldots,H_{r}$ are hyperplanes in $\mathbb{R}^{d}$ such that (1) the $r$-fold intersection of the hyperplanes is a flat of codimension $k+1$ and (2) any $k+1$ of the $r$ hyperplanes are independent (i.e., their normal vectors are linearly independent). Taking indices modulo $r$, for each $j\in[r]$ let $A_{j}$ be the affine space in $\mathbb{R}^{d}$ of codimension $k$ given by the $k$-fold intersection $A_{j}=\cap_{i=j+1}^{j+k}H_{i}$ of the hyperplanes and let $B_{j}=H_{j}^{+}\cap A_{j}$ be one of the (closed) half-flats of $A_{j}$ determined by a half-space $H_{j}^{+}$ of $H_{j}$. We call the union $F_{r}=\cup_{j=1}^{r}B_{j}$ of these half-flats an $r$-fan of codimenison $k$. The boundary of each half-flat is the common intersection $C=\cap_{i=1}^{r}H_{i}$ of the $r$ hyerplanes, which we call the center of the fan. In particular, an $r$-fan of codimension one is the union of $r$ half-hyperplanes centered about an affine space of codimension two. Equipartitions by the regions determined by such $r$-fans have been previously considered (see, e.g., [2, 3, 30]). We say that an $r$-fan of codimension one is regular if the dihedral angle between each of its successive half-hyperplanes is equal to $2\pi/r$.
@@
We have the following equidsitribution result for $r$-fans of codimension $r-2$. Recall that a set of points in $\mathbb{R}^{d}$ is affinely spanning if it is not contained in a hyperplane.
@@
As with Theorem 1.1, the $r=2$ case of Theorem 1.3 is nearly sharp with respect to complex dimension (see Remark 5.4 below). As a regular $2$-fan in $\mathbb{R}^{d}$ is a hyperplane, the $r=2$ case of Theorem  1.3 is closely aligned with the ham sandwich theorem: for any $m$-coloring $X=X_{1}\sqcup\cdots\sqcup X_{m}$ of a set $X$ of $n$ complex affinely spanning points in $\mathbb{C}^{n-d-1}\cong\mathbb{R}^{2n-2d-2}$, there is a real hyperplane $H$ in $\mathbb{C}^{n-d-1}$ containing $X$ and a complex hyperplane $H_{\mathbb{C}}$ containing $H$ such that $H_{\mathbb{C}}$, when viewed as a real hyperplane in $H$, equipartitions each of the $X_{i}$.
@@ 1.2. Piercing Distributions
The $r=2$ case of Theorem 1.4(b) gives an analogue of Dolnikov’s theorem for complex instead of real hyperplanes: If $X$ is a set of $n$ complex affinely spanning points in $\mathbb{C}^{n-d-1}$ and $\mathcal{F}_{1},\ldots,\mathcal{F}_{m}$ are non-empty intersecting families of subsets of $X$, then there exists a real hyperplane $H$ in $\mathbb{C}^{n-d-1}$ which contains $X$ and a complex hyperplane $H_{\mathbb{C}}$ contained in $H$ which intersects the convex hull of every member of each $\mathcal{F}_{i}$.
@@ 1.3. Proof Scheme
Recall that Tverberg’s theorem [32] asserts that any set of $n\geq(r-1)(d+1)+1$ points in $\mathbb{R}^{d}$ can be partitioned into $r$ pairwise disjoint sets whose convex hulls have non-empty $r$-fold intersection. The $r=2$ case is Radon’s theorem. For prime powers, one has a topological extension of Tverberg theorem [22, 33] which asserts that any continuous map $f\colon\Delta_{n-1}\rightarrow\mathbb{R}^{d}$ from the $(n-1)$-dimenisional simplex to $\mathbb{R}^{d}$ has an $r$-Tverberg tuple – that is, an $r$-tuple $(\sigma_{1},\ldots,\sigma_{r})$ of non-empty pairwise disjoint faces of $\Delta_{n-1}$ such that $f(\sigma_{1})\cap\cdots\cap f(\sigma_{r})\neq\emptyset$. When $r$ is a prime power, Tverberg’s theorem is recovered by letting $f\colon\Delta_{n-1}\rightarrow\mathbb{R}^{d}$ be a linear map, that is, one which sends each convex combination of vertices of the simplex to the corresponding convex combination of the image of the vertices. In what follows, we identify the vertex set of $\Delta_{n-1}$ with $[n]=\{1,\ldots,n\}$ and we let $\operatorname{\mathrm{Vert}}(\sigma)\subseteq[n]$ denote the vertex set of any non-empty face $\sigma$ of $\Delta_{n-1}$.
@@ 2. Real and Complex Gale Duality
For our purposes, the essential property of the Gale transform is that it gives a correspondence between any affine dependency among the $a_{1},\ldots,a_{n}$ and an allocation of the $g_{1},\ldots,g_{n}$ with respect to a codimenison one linear subspace. More precisely, let $\langle u,w\rangle=\sum_{i=1}^{n}u_{i}\overline{w}_{i}$ denote the standard Hermitian inner product of two vectors $u$ and $w$ in $\mathbb{K}^{n}$. One then has the following.
@@ 3.1. Real Linear Fans
For part (b), suppose that $F_{r}$ a linear $r$-fan in $\mathbb{R}^{n-d-1}$ of codimenison $r-2$ determined by $H_{1}=\langle\alpha_{1}\rangle^{\perp},\ldots,H_{r}=\langle\alpha_{r}\rangle^{\perp}$ in $\mathbb{R}^{n-d-1}$. By assumption, $\alpha_{1},\ldots,\alpha_{r}$ are linearly dependent, so $\sum_{j=1}^{r}\mu_{j}\alpha_{j}=0$ for some scalars $\mu_{1},\ldots,\mu_{r}$, not all of which are zero. Since any $r-1$ of the $\alpha_{j}$ are linearly independent, $\mu_{j}\neq 0$ for all $j\in[r]$. Thus by scaling both $\alpha_{j}$ and $\beta_{j}$ by $\mu_{j}$ we may assume that $\sum_{j=1}^{r}\alpha_{j}=0$. Without loss of generality, we may let $H^{+}_{j}=\{x\in\mathbb{R}^{n-d-1}\mid\langle\alpha_{j},x\rangle\geq 0\}$ for all $j\in[r]$.
@@ 4. Distributions by a Single Fan
We begin by proving our equdistribution results for a single fan, which are a special case of Theorem 1.4.
@@
In the complex setting, we have a complex regular $r$-fan $F_{r}=\cup_{j=1}^{r}B_{j}$ in $\mathbb{C}^{n-d}$ centered about a linear complex hyperplane $H_{\mathbb{C}}$. As above, we let $H_{\mathbb{C}}^{\prime}=H_{\mathbb{C}}\cap(\mathbb{C}^{n-d-1}\times\{1\})$, $B_{j}^{\prime}=B_{j}\cap(\mathbb{C}^{n-d-1}\times\{1\})$ for each $j\in[r]$, and we let $H^{\prime\prime}_{\mathbb{C}}$ and $B^{\prime\prime}_{j}$ be their corresponding projections onto $\mathbb{C}^{n-d-1}$. We claim that $F_{r}^{\prime\prime}=\cup_{j=1}^{r}B^{\prime\prime}_{j}$ is a regular $r$-fan in $\mathbb{C}^{n-d-1}$ centered about the complex hyerplane $H^{\prime\prime}_{\mathbb{C}}$. To verify this, we have $H_{\mathbb{C}}=\langle\alpha\rangle^{\perp}$, where $\alpha=(\beta,\gamma)\in\mathbb{C}^{n-d-1}\times\mathbb{C}$ is non-zero. Thus $H^{\prime\prime}_{\mathbb{C}}=\{z\in\mathbb{C}^{n-d-1}\mid\langle\beta,z\rangle=-\gamma\}$ and $B^{\prime}_{j}=\{(z\in\mathbb{C}^{n-d-1}\mid\langle\beta,z\rangle+\gamma=t\omega_{r}^{j}\,\,\text{for some}\,\,t\geq 0\}$. Thus $H_{\mathbb{C}}^{\prime\prime}$ and $F_{r}^{\prime\prime}$ are as claimed provided $\beta\neq 0$. But $g_{n+1}=(-x,-n)\in H_{\mathbb{C}}$ where $x=\sum_{i=1}^{n}g_{i}$. Thus $\langle(x,-n),(\beta,\gamma)\rangle=0$ and so $\langle\beta,x\rangle=n\gamma$. But if $\beta=0$, then $\gamma=0$ and therefore $\alpha=0$ as well. Thus $F^{\prime\prime}_{r}$ is indeed a regular $r$-fan centered about the complex hyperplane $H_{\mathbb{C}}^{\prime\prime}$.
@@
We conclude this section by proving our colorful distribution result. This follows the same procedure as before, except we substitue Sarkaria’s theorem with the Optimal Colored Tverberg theorem of  [8] below.
@@ 5.1. Typical Point Sets
    Let $r=2$, let $k\in\{0,1\}$, and let $n=2(d+1)+m+k$. If $\ell>2+\frac{k}{2}$, then there exists a set $X$ of $n+\ell$ complex affinely spanning points in $\mathbb{C}^{n-d-1}$ and an $m$-coloring of $X$ which cannot be equdistributed by any complex regular $2$-fan in $\mathbb{C}^{n-d-1}$.
@@
We now show that there exists an $m$-coloring of $X=\{x_{1},\ldots,x_{n+\ell}\}$ which cannot be equidstributed by any $r$-fan in $\mathbb{R}^{n-d-1}$ of codimension $r-2$. Namely, let $X=X_{1}\sqcup\cdots\sqcup X_{m}$ be an $m$-coloring such that $|X_{i}|=r-1$ for all $i\geq 2$ and $x_{n+\ell}=0$ is in $X_{m}$. Now suppose that $F_{r}=\cup_{j=1}^{r}B_{j}$ is an $r$-fan of codimension $r-2$ which equidistributes this coloring. As $|\operatorname{\mathrm{Int}}(B_{j})\cap X_{i}|\leq|X_{i}|/r$ for all $i\in[m]$ and $j\in[r]$, all the points of $X\setminus X_{1}$ must lie on the center $C$ of $F_{r}$. In particular, $F_{r}$ is linear. Now let $f\colon\Delta_{n+\ell-1}\rightarrow\mathbb{R}^{d+\ell}$ be the linear map determined by the $a_{i}^{\prime}$ and let $[n+\ell]=C_{1}\sqcup\cdots\sqcup C_{m}$ be the $m$-coloring of the vertex set of $\Delta_{n+\ell-1}$ corresponding to the $m$-coloring of $X$. Applying part (b) of Lemma 3.1, there is a proper Tverberg $r$-tuple $(\sigma_{1},\ldots,\sigma_{r})$ for $f$ such that each $I_{j}:=\operatorname{\mathrm{Vert}}(\sigma_{j})$ is in bijective correspondence with $X\cap\operatorname{\mathrm{Int}}(B_{j})$ and $[n+\ell-1]\setminus\sqcup_{j=1}^{r}I_{j}$ is in bijective correspondence with $C\cap X$. Thus $|C_{i}\cap\operatorname{\mathrm{Vert}}(\sigma)|=|X_{i}\cap\operatorname{\mathrm{Int}}(B_{j})|$ for all $j\in[r]$ and all $i\in[m]$. Thus each $I_{j}$ is contained in the color class $C_{1}\subset[n+\ell-1]$, which consists of $n-(m-1)(r-1)=(r-1)(d+2)+k+1$ points. Letting $\Delta_{C_{1}}$ be the simplex generated by $C_{1}$, we have that the image of the restriction $f\colon\Delta_{C_{1}}\rightarrow\mathbb{R}^{d+\ell}$ of $f$ has image in $\mathbb{R}^{d+\ell-1}\times\{1\}$. Since the points $a_{1}^{\prime}=(a_{1},1),\ldots,a_{n+\ell-1}^{\prime}=(a_{n+\ell-1},1)$ are in strong general position in $\mathbb{R}^{d+\ell-1}\times\{1\}$ and $k<(r-1)(\ell-2)$, we have
@@
As in the real case, we may denote by $n_{\mathbb{C}}(r,m,d)$ the maximum number $n$ such that any $m$-coloring of any set of $n$ complex affinely spanning points in $\mathbb{C}^{d}$ can be equidistributed by a complex regular $r$-fan in $\mathbb{C}^{d}$. Letting $d=s+t+m+1$ with $s\geq 1$ and $t\in\{0,1\}$ and setting $n=d+s+1$, Theorem 1.3 and Proposition 5.3 likewise give the bound $n\leq n_{\mathbb{C}}(2,m,d)\leq n+2$ for complex regular $2$-fans.

### v2

@@ Abstract.
Equipartition theory, beginning with the classical ham sandwich theorem, seeks the fair division of finite point sets in $\mathbb{R}^{d}$ by the full-dimensional regions determined by a prescribed geometric dissection of $\mathbb{R}^{d}$. Here we examine equidistributions of finite point sets in $\mathbb{R}^{d}$ by prescribed low dimensional subsets. Our main result states that if $r\geq 3$ is a prime power, then for any $m$-coloring of a sufficiently small point set $X$ in $\mathbb{R}^{d}$, there exists an $r$-fan in $\mathbb{R}^{d}$ – that is, the union of $r$ “half-flats” of codimension $r-2$ centered about a common $(r-1)$-codimensional affine subspace – which captures all the points of $X$ in such a way that each half-flat contains at most an $r$-th of the points from each color class. The number of points in $\mathbb{R}^{d}$ we require for this is essentially tight when $m\geq 2$. Additionally, we extend our equidistribution results to ”piercing” distributions in a similar fashion to Dolnikov’s hyperplane transversal generalization of the ham sandwich theorem. By analogy with recent work of Frick et al., our results are obtained by applying Gale duality to linear cases of topological Tverberg-type theorems. Finally, we extend our distribution results to multiple $r$-fans after establishing a multiple intersection version of a topological Tverberg-type theorem due to Sarkaria.
@@ 1.1. From Equipartitions to Equidistributions of Colored Point Sets
Instead of equipartitions of finite point sets in $\mathbb{R}^{d}$ by full dimensional regions, here we seek their equidistribution by small-dimensional subsets of $\mathbb{R}^{d}$ arising from families of hyperplane with prescribed intersections. Specifically, suppose that $1\leq k\leq\min\{d-1,r-1\}$ and that $H_{1},\ldots,H_{r}$ are hyperplanes in $\mathbb{R}^{d}$ such that (1) the $r$-fold intersection of the hyperplanes is a flat of codimension $k+1$ and (2) any $k+1$ of the $r$ hyperplanes are independent (i.e., their normal vectors are linearly independent). Taking indices modulo $r$, for each $j\in[r]$ let $A_{j}$ be the affine space in $\mathbb{R}^{d}$ of codimension $k$ given by the $k$-fold intersection $A_{j}=\cap_{i=j+1}^{j+k}H_{i}$ of the hyperplanes and let $B_{j}=H_{j}^{+}\cap A_{j}$ be one of the (closed) half-flats of $A_{j}$ determined by a half-space $H_{j}^{+}$ of $H_{j}$. We call the union $F_{r}=\cup_{j=1}^{r}B_{j}$ of these half-flats an $r$-fan of codimension $k$. The boundary of each half-flat is the common intersection $C=\cap_{i=1}^{r}H_{i}$ of the $r$ hyperplanes, which we call the center of the fan. In particular, an $r$-fan of codimension one is the union of $r$ half-hyperplanes centered about an affine space of codimension two. Equipartitions by the regions determined by such $r$-fans have been previously considered (see, e.g., [2, 3, 30]). We say that an $r$-fan of codimension one is regular if the dihedral angle between each of its successive half-hyperplanes is equal to $2\pi/r$.
@@
We have the following equidistribution result for $r$-fans of codimension $r-2$. Recall that a set of points in $\mathbb{R}^{d}$ is affinely spanning if it is not contained in a hyperplane.
@@
As a regular $2$-fan in $\mathbb{R}^{d}$ is a hyperplane, the $m=2$ case of Theorem  1.3 is closely aligned with the ham sandwich theorem. Namely, let $d=s+m+1+t$, where $s\geq 1$ and $t\in\{0,1\}$. Then for any $m$-coloring $X=X_{1}\sqcup\cdots\sqcup X_{m}$ of a set $X$ of $n:=d+s+1$ complex affinely spanning points in $\mathbb{C}^{d}$, there is a real hyperplane $H$ in $\mathbb{C}^{d}$ containing $X$ and a complex hyperplane $H_{\mathbb{C}}$ contained in $H$ such that $H_{\mathbb{C}}$, when viewed as a real hyperplane in $H$, equipartitions each of the $X_{i}$. The number of points in $\mathbb{C}^{d}$ required for this is exceeds the optimum number by at most $2$ for any $m\geq 1$ (see Remark 5.4 below).
@@ 1.2. Piercing Distributions
As an example of Theorem 1.4(a), letting $r=4$ and $d=1$ shows that for any affinely spanning set $X$ of $10$ points in $\mathbb{R}^{8}$, there is a $6$-dimensional $4$-fan in $\mathbb{R}^{8}$ which contains $X$ so that none of the three element subsets of $X$ is contained in the interior of any of the fan’s half-flats. If $X$ is typical, then the interiors of the fan’s half-flats must share at least $7$ of the $10$ points.
The $r=2$ case of Theorem 1.4(b) gives an analogue of Dolnikov’s theorem for complex instead of real hyperplanes: If $d=s+m+1+t$ with $s\geq 1$ and $t\in\{0,1\}$, let $X$ be an affinely spanning set of $n:=d+s+1$ points in $\mathbb{C}^{d}$ and let $\mathcal{F}_{1},\ldots,\mathcal{F}_{m}$ be non-empty intersecting families of subsets of $X$. Then exists a real hyperplane $H$ in $\mathbb{C}^{d}$ which contains $X$ and a complex hyperplane $H_{\mathbb{C}}$ contained in $H$ such that $H_{\mathbb{C}}$ intersects the convex hull of every member of each $\mathcal{F}_{i}$.
@@ 1.3. Proof Scheme
Recall that Tverberg’s theorem [32] asserts that any set of $n\geq(r-1)(d+1)+1$ points in $\mathbb{R}^{d}$ can be partitioned into $r$ pairwise disjoint sets whose convex hulls have non-empty $r$-fold intersection. The $r=2$ case is Radon’s theorem. For prime powers, one has a topological extension of Tverberg theorem [22, 33] which asserts that any continuous map $f\colon\Delta_{n-1}\rightarrow\mathbb{R}^{d}$ from the $(n-1)$-dimensional simplex to $\mathbb{R}^{d}$ has an $r$-Tverberg tuple – that is, an $r$-tuple $(\sigma_{1},\ldots,\sigma_{r})$ of non-empty pairwise disjoint faces of $\Delta_{n-1}$ such that $f(\sigma_{1})\cap\cdots\cap f(\sigma_{r})\neq\emptyset$. When $r$ is a prime power, Tverberg’s theorem is recovered by letting $f\colon\Delta_{n-1}\rightarrow\mathbb{R}^{d}$ be a linear map, that is, one which sends each convex combination of vertices of the simplex to the corresponding convex combination of the image of the vertices. In what follows, we identify the vertex set of $\Delta_{n-1}$ with $[n]=\{1,\ldots,n\}$ and we let $\operatorname{\mathrm{Vert}}(\sigma)\subseteq[n]$ denote the vertex set of any non-empty face $\sigma$ of $\Delta_{n-1}$.
@@
As an example of Theorem 1.6, again let $r=4$ and $d=1$. Then for any partition of a set of $8$ affinely spanning points in $\mathbb{R}^{6}$ into two color classes of size $4$ each, there is a $4$-dimensional $4$-fan in $\mathbb{R}^{6}$ such that the interior of each half-flat contains at most one point from each color class. Moreover, if $X$ is typical then the interiors of the fan’s half-flats must share at least $7$ of the $8$ points.
@@ 2. Real and Complex Gale Duality
For our purposes, the essential property of the Gale transform is that it gives a correspondence between any affine dependency among the $a_{1},\ldots,a_{n}$ and an allocation of the $g_{1},\ldots,g_{n}$ with respect to a codimension one linear subspace. More precisely, let $\langle u,w\rangle=\sum_{i=1}^{n}u_{i}\overline{w}_{i}$ denote the standard Hermitian inner product of two vectors $u$ and $w$ in $\mathbb{K}^{n}$. One then has the following.
@@ 3.1. Real Linear Fans
For part (b), suppose that $F_{r}$ a linear $r$-fan in $\mathbb{R}^{n-d-1}$ of codimension $r-2$ determined by $H_{1}=\langle\alpha_{1}\rangle^{\perp},\ldots,H_{r}=\langle\alpha_{r}\rangle^{\perp}$ in $\mathbb{R}^{n-d-1}$. By assumption, $\alpha_{1},\ldots,\alpha_{r}$ are linearly dependent, so $\sum_{j=1}^{r}\mu_{j}\alpha_{j}=0$ for some scalars $\mu_{1},\ldots,\mu_{r}$, not all of which are zero. Since any $r-1$ of the $\alpha_{j}$ are linearly independent, $\mu_{j}\neq 0$ for all $j\in[r]$. Thus by scaling both $\alpha_{j}$ and $\beta_{j}$ by $\mu_{j}$ we may assume that $\sum_{j=1}^{r}\alpha_{j}=0$. Without loss of generality, we may let $H^{+}_{j}=\{x\in\mathbb{R}^{n-d-1}\mid\langle\alpha_{j},x\rangle\geq 0\}$ for all $j\in[r]$.
@@ 4. Distributions by a Single Fan
We begin by proving our equidistribution results for a single fan, which are a special case of Theorem 1.4.
@@
In the complex setting, we have a complex regular $r$-fan $F_{r}=\cup_{j=1}^{r}B_{j}$ in $\mathbb{C}^{n-d}$ centered about a linear complex hyperplane $H_{\mathbb{C}}$. As above, we let $H_{\mathbb{C}}^{\prime}=H_{\mathbb{C}}\cap(\mathbb{C}^{n-d-1}\times\{1\})$, $B_{j}^{\prime}=B_{j}\cap(\mathbb{C}^{n-d-1}\times\{1\})$ for each $j\in[r]$, and we let $H^{\prime\prime}_{\mathbb{C}}$ and $B^{\prime\prime}_{j}$ be their corresponding projections onto $\mathbb{C}^{n-d-1}$. We claim that $F_{r}^{\prime\prime}=\cup_{j=1}^{r}B^{\prime\prime}_{j}$ is a regular $r$-fan in $\mathbb{C}^{n-d-1}$ centered about the complex hyperplane $H^{\prime\prime}_{\mathbb{C}}$. To verify this, we have $H_{\mathbb{C}}=\langle\alpha\rangle^{\perp}$, where $\alpha=(\beta,\gamma)\in\mathbb{C}^{n-d-1}\times\mathbb{C}$ is non-zero. Thus $H^{\prime\prime}_{\mathbb{C}}=\{z\in\mathbb{C}^{n-d-1}\mid\langle\beta,z\rangle=-\gamma\}$ and $B^{\prime}_{j}=\{(z\in\mathbb{C}^{n-d-1}\mid\langle\beta,z\rangle+\gamma=t\omega_{r}^{j}\,\,\text{for some}\,\,t\geq 0\}$. Thus $H_{\mathbb{C}}^{\prime\prime}$ and $F_{r}^{\prime\prime}$ are as claimed provided $\beta\neq 0$. But $g_{n+1}=(-x,-n)\in H_{\mathbb{C}}$ where $x=\sum_{i=1}^{n}g_{i}$. Thus $\langle(x,-n),(\beta,\gamma)\rangle=0$ and so $\langle\beta,x\rangle=n\gamma$. But if $\beta=0$, then $\gamma=0$ and therefore $\alpha=0$ as well. Thus $F^{\prime\prime}_{r}$ is indeed a regular $r$-fan centered about the complex hyperplane $H_{\mathbb{C}}^{\prime\prime}$.
@@
We conclude this section by proving our colorful distribution result. This follows the same procedure as before, except we substitute Sarkaria’s theorem with the Optimal Colored Tverberg theorem of  [8] below.
@@ 5.1. Typical Point Sets
    Let $r=2$, let $k\in\{0,1\}$, and let $n=2(d+1)+m+k$. If $\ell>2+\frac{k}{2}$, then there exists a set $X$ of $n+\ell$ complex affinely spanning points in $\mathbb{C}^{n-d-1}$ and an $m$-coloring of $X$ which cannot be equidistributed by any complex regular $2$-fan in $\mathbb{C}^{n-d-1}$.
@@
We now show that there exists an $m$-coloring of $X=\{x_{1},\ldots,x_{n+\ell}\}$ which cannot be equidistributed by any $r$-fan in $\mathbb{R}^{n-d-1}$ of codimension $r-2$. Namely, let $X=X_{1}\sqcup\cdots\sqcup X_{m}$ be an $m$-coloring such that $|X_{i}|=r-1$ for all $i\geq 2$ and $x_{n+\ell}=0$ is in $X_{m}$. Now suppose that $F_{r}=\cup_{j=1}^{r}B_{j}$ is an $r$-fan of codimension $r-2$ which equidistributes this coloring. As $|\operatorname{\mathrm{Int}}(B_{j})\cap X_{i}|\leq|X_{i}|/r$ for all $i\in[m]$ and $j\in[r]$, all the points of $X\setminus X_{1}$ must lie on the center $C$ of $F_{r}$. In particular, $F_{r}$ is linear. Now let $f\colon\Delta_{n+\ell-1}\rightarrow\mathbb{R}^{d+\ell}$ be the linear map determined by the $a_{i}^{\prime}$ and let $[n+\ell]=C_{1}\sqcup\cdots\sqcup C_{m}$ be the $m$-coloring of the vertex set of $\Delta_{n+\ell-1}$ corresponding to the $m$-coloring of $X$. Applying part (b) of Lemma 3.1, there is a proper Tverberg $r$-tuple $(\sigma_{1},\ldots,\sigma_{r})$ for $f$ such that each $I_{j}:=\operatorname{\mathrm{Vert}}(\sigma_{j})$ is in bijective correspondence with $X\cap\operatorname{\mathrm{Int}}(B_{j})$ and $[n+\ell-1]\setminus\sqcup_{j=1}^{r}I_{j}$ is in bijective correspondence with $C\cap X$. Thus $|C_{i}\cap\operatorname{\mathrm{Vert}}(\sigma)|=|X_{i}\cap\operatorname{\mathrm{Int}}(B_{j})|$ for all $j\in[r]$ and all $i\in[m]$. Thus each $I_{j}$ is contained in the color class $C_{1}\subset[n+\ell-1]$, which consists of $n-(m-1)(r-1)=(r-1)(d+2)+k+1$ points. Letting $\Delta_{C_{1}}$ be the simplex generated by $C_{1}$, we have that the image of the restriction $f\colon\Delta_{C_{1}}\rightarrow\mathbb{R}^{d+\ell}$ of $f$ has image in $\mathbb{R}^{d+\ell-1}\times\{1\}$. Since the points $a_{1}^{\prime}=(a_{1},1),\ldots,a_{n+\ell-1}^{\prime}=(a_{n+\ell-1},1)$ are in strong general position in $\mathbb{R}^{d+\ell-1}\times\{1\}$ and $k<(r-1)(\ell-2)$, we have
@@
As in the real case, we denote by $n_{\mathbb{C}}(r,m,d)$ the maximum number $n$ such that any $m$-coloring of any set of $n$ complex affinely spanning points in $\mathbb{C}^{d}$ can be equidistributed by a complex regular $r$-fan in $\mathbb{C}^{d}$. Letting $d=s+t+m+1$ with $s\geq 1$ and $t\in\{0,1\}$ and setting $n=d+s+1$, Theorem 1.3 together with Proposition 5.3 likewise give the bound $n\leq n_{\mathbb{C}}(2,m,d)\leq n+2$ for complex regular $2$-fans when $m\geq 2$. When $m=1$, one has $n=2s+t+3$ points in $\mathbb{C}^{d}\cong\mathbb{R}^{2d}$. Considering points which are in ordinary general position in $\mathbb{C}^{d}$, no more than $42d=2s+2t+$ 4 points can lie on any real hyperplane. Thus $n\leq n_{\mathbb{C}}(2,1,d)\leq n+2$ as well.